OpenStudy (anonymous):

solves this problem: square root of (x-1)=10

OpenStudy (anonymous):

can someone help me step by step....???

OpenStudy (anonymous):

\[\sqrt{x-10} = 10 \Rightarrow (\sqrt{x-10})^{2} = 10^{2} \Rightarrow x-10 = 100 \Rightarrow x = 110\]

OpenStudy (anonymous):

But you actually made a lil mistake, buddy :P

OpenStudy (anonymous):

first square both sides, linhphan. So you get: (x-1) = 100 Then you add the 1 (joe accidentally put down 10) so x = 101

OpenStudy (anonymous):

whoops, my bad >.< thanks for catching that :)

OpenStudy (anonymous):

No problemo. ^_^

OpenStudy (anonymous):

hiihi thanks:)...